The difference between two numbers is 16. Five times the smaller is the same aas 8 less that twice the larger. Find the numbers. I need this to be in a system of two equations in two variables.
Answer by checkley75(3666)

X-Y=16 X=Y+16 NOW SUBSTITUTE (Y+16) FOR X IN THE OTHER EQUATION
5X-8=2Y THUS
5(Y+16)-8=2Y
5Y+80-8=2Y
5Y-2Y=+8-80
3Y=-72
Y=-72/3
Y=-24 THUS
X-(-24)=16
X+24=16
X=16-24
X=-8
PROOF
5*-8-8=2*-24
-40-8=-48
-48=-48